OpenStudy (anonymous):

can you give the steps for 4 8/10 divided by 4/5

OpenStudy (anonymous):

1) Change 4 and 8/10ths into an improper fraction. 2) Invert the second fraction (flip the numberator and denominator. 3) Multiply the two fractions together. \[\frac{48}{10}*\frac{5}{4}=\frac{12}{2}\] <<-- Notice that I cross canceled instead. I'll let you reduce the fraction:-)